# SMMILE: An Expert-Driven Benchmark for Multimodal Medical In-Context Learning

### v1

@@ 3.2 Benchmarking MLLMs with SMMILE
LLM-as-a-Judge EM MCQA
0-shot ICL 0-shot ICL 0-shot ICL
Majority $26.30\pm\scriptstyle{3.88}$ $27.26\pm\scriptstyle{4.03}$ $24.15\pm\scriptstyle{3.70}$
Random $27.86\pm\scriptstyle{4.43}$ $23.16\pm\scriptstyle{3.88}$ $36.30\pm\scriptstyle{4.66}$
Text only $5.32\pm\scriptstyle{2.27}$ $16.53\pm\scriptstyle{3.94}$ $0.00\pm\scriptstyle{0.00}$ $5.22\pm\scriptstyle{1.70}$ $38.62\pm\scriptstyle{4.61}$ $28.03\pm\scriptstyle{4.28}$
Claude 3.7 Sonnet $\mathbf{37.18}\pm\scriptstyle{4.39}$ $36.17\pm\scriptstyle{4.44}$ $0.00\pm\scriptstyle{0.00}$ $2.63\pm\scriptstyle{1.67}$ $\mathbf{56.10}\pm\scriptstyle{4.31}$ $42.01\pm\scriptstyle{4.83}$
GPT-4o $32.56\pm\scriptstyle{4.60}$ $\mathbf{49.88}\pm\scriptstyle{4.69}$ $0.00\pm\scriptstyle{0.00}$ $8.94\pm\scriptstyle{2.54}$ $49.74\pm\scriptstyle{4.48}$ $\mathbf{58.85}\pm\scriptstyle{4.62}$
Llama-3.2-Vision-90B $31.84\pm\scriptstyle{4.38}$ $40.66\pm\scriptstyle{4.99}$ $0.00\pm\scriptstyle{0.00}$ $30.53\pm\scriptstyle{4.07}$ $55.04\pm\scriptstyle{4.93}$ $30.30\pm\scriptstyle{5.20}$
LLaVA-v1.5-7B $14.61\pm\scriptstyle{3.57}$ $18.72\pm\scriptstyle{3.40}$ $0.00\pm\scriptstyle{0.00}$ $16.37\pm\scriptstyle{3.32}$ $40.34\pm\scriptstyle{5.35}$ $22.30\pm\scriptstyle{3.92}$
LLaVA-v1.5-13B $19.58\pm\scriptstyle{3.64}$ $20.91\pm\scriptstyle{3.49}$ $0.00\pm\scriptstyle{0.00}$ $19.54\pm\scriptstyle{3.95}$ $38.96\pm\scriptstyle{4.83}$ $24.92\pm\scriptstyle{4.25}$
LLaVA-NeXT-7B $11.42\pm\scriptstyle{3.04}$ $23.66\pm\scriptstyle{3.90}$ $0.00\pm\scriptstyle{0.00}$ $2.69\pm\scriptstyle{1.31}$ $38.11\pm\scriptstyle{4.26}$ $29.01\pm\scriptstyle{4.05}$
LLaVA-OneVision-0.5B $18.26\pm\scriptstyle{3.93}$ $21.63\pm\scriptstyle{4.00}$ $2.61\pm\scriptstyle{1.59}$ $13.46\pm\scriptstyle{3.11}$ $44.03\pm\scriptstyle{4.47}$ $32.11\pm\scriptstyle{4.44}$
LLaVA-OneVision-7B $16.41\pm\scriptstyle{3.65}$ $24.25\pm\scriptstyle{3.81}$ $3.44\pm\scriptstyle{2.00}$ $22.43\pm\scriptstyle{4.33}$ $40.15\pm\scriptstyle{4.59}$ $27.17\pm\scriptstyle{4.28}$
LLaVA-Med-7B∗∗ $21.65\pm\scriptstyle{4.18}$ $10.19\pm\scriptstyle{3.06}$ $0.00\pm\scriptstyle{0.00}$ $0.00\pm\scriptstyle{0.00}$ $0.00\pm\scriptstyle{0.00}$ $0.00\pm\scriptstyle{0.00}$
MedGemma-4B-Multimodal $27.73\pm\scriptstyle{4.70}$ $36.86\pm\scriptstyle{4.81}$ $0.00\pm\scriptstyle{0.00}$ $12.14\pm\scriptstyle{3.07}$ $41.21\pm\scriptstyle{5.15}$ $40.67\pm\scriptstyle{4.93}$
MedVLM-R1 $25.20\pm\scriptstyle{4.09}$ $26.74\pm\scriptstyle{4.44}$ $\mathbf{3.65}\pm\scriptstyle{1.50}$ $15.26\pm\scriptstyle{2.91}$ $36.54\pm\scriptstyle{4.63}$ $26.22\pm\scriptstyle{4.31}$
Qwen2.5-VL-3B $27.62\pm\scriptstyle{4.26}$ $33.58\pm\scriptstyle{4.09}$ $0.00\pm\scriptstyle{0.00}$ $26.18\pm\scriptstyle{4.52}$ $37.58\pm\scriptstyle{4.41}$ $27.35\pm\scriptstyle{4.23}$
Qwen2.5-VL-7B $17.90\pm\scriptstyle{3.20}$ $29.58\pm\scriptstyle{4.63}$ $0.00\pm\scriptstyle{0.00}$ $22.45\pm\scriptstyle{3.81}$ $38.24\pm\scriptstyle{4.75}$ $45.01\pm\scriptstyle{4.39}$
Qwen2.5-VL-32B $25.27\pm\scriptstyle{3.86}$ $41.79\pm\scriptstyle{4.73}$ $0.00\pm\scriptstyle{0.00}$ $\mathbf{31.84}\pm\scriptstyle{4.37}$ $51.76\pm\scriptstyle{4.46}$ $49.97\pm\scriptstyle{5.07}$
Qwen2.5-VL-72B $29.90\pm\scriptstyle{4.08}$ $42.59\pm\scriptstyle{4.55}$ $0.00\pm\scriptstyle{0.00}$ $15.71\pm\scriptstyle{3.30}$ $52.33\pm\scriptstyle{4.58}$ $54.71\pm\scriptstyle{4.89}$
@@
  • There are notable changes in model rankings. Unlike Table 1 where GPT-4o was dominant, Qwen2.5-VL-72B now takes over the lead (53.8% LLM-as-a-Judge accuracy, 63.2% on ICL-MCQA).
  • Broader benefits from in-context learning are visible. Broader benefits from in-context learning are visible. The largest relative performance improvements are observed when in-context examples are presented to models, such as: LLaVA-v1.5-7B (99.4% relative improvement in LLM-as-a-Judge, 12.31% → 24.55%), Qwen2.5-VL-7B (94.4% relative improvement in LLM-as-a-Judge, 21.10% → 41.01%), and Qwen2.5-VL-3B (79.9% relative improvement in LLM-as-a-Judge, 17.53% → 31.54%). In the open-ended setting, all models (with the notable exception of LLaVA-Med-7B) demonstrate higher ICL performance than zero-shot performance when evaluated with LLM-as-a-Judge, with an average relative improvement of 44.7%.
@@
Majority - $17.70\pm\scriptstyle{1.19}$ $17.59\pm\scriptstyle{1.19}$ $16.65\pm\scriptstyle{1.13}$
Random - $25.35\pm\scriptstyle{1.34}$ $25.35\pm\scriptstyle{1.32}$ $33.77\pm\scriptstyle{1.46}$
@@
GPT-4o $\mathbf{38.41}\pm\scriptstyle{1.51}$ $46.45\pm\scriptstyle{1.54}$ $0.00\pm\scriptstyle{0.00}$ $7.79\pm\scriptstyle{0.81}$ $56.70\pm\scriptstyle{1.48}$ $55.76\pm\scriptstyle{1.56}$
@@
LLaVA-Med-7B∗∗ $24.84\pm\scriptstyle{1.31}$ $4.62\pm\scriptstyle{0.64}$ $0.00\pm\scriptstyle{0.00}$ $0.19\pm\scriptstyle{0.14}$ $0.29\pm\scriptstyle{0.17}$ $0.00\pm\scriptstyle{0.00}$
MedGemma-4B-Multimodal $24.72\pm\scriptstyle{1.35}$ $38.66\pm\scriptstyle{1.54}$ $0.00\pm\scriptstyle{0.00}$ $13.30\pm\scriptstyle{1.06}$ $40.36\pm\scriptstyle{1.55}$ $44.78\pm\scriptstyle{1.52}$
@@ 3.3 Fine-Grained Analysis
We now perform a fine-grained breakdown of MLLM performance across the SMMILE benchmark. We specifically focus on five reproducible MLLMs for this analysis: LLaVA-OneVision-0.5B, LLaVA-Med-7B, LLaVA-v1.5-13B, Qwen2.5-VL-32B, and Qwen2.5-VL-72B. We first evaluate MLLM performance stratified by answer format. Each ground-truth answer in the SMMILE dataset was annotated by an expert with one of the following four categorical labels: "multiple choice", "free response", "binary (yes/no)", or "numerical". As shown in Figure 4 (Panel A), MLLMs display substantial variations in performance across the four categories, with all five evaluated models demonstrating the strongest performance on binary (yes/no) answers. Notably, we find that all evaluated models fail to correctly answer questions with numerical answers, which is a critical limitation since the ability to provide quantitative responses is vital for effective decision-making in medical settings. This finding is corroborated by analysis in Figure 4 (Panel B), which again finds that all evaluated MLLMs struggle when quantitative reasoning is necessary to answer a medical question.
@@
In Figure 4 (Panel C), we stratify SMMILE by the number of in-context examples included with each question. We report the effect of the number of ICL examples on MLLM performance, and we also include results from zero-shot evaluations (where no examples are provided to the model). We find that for all evaluated models, providing two ICL examples leads to substantial improvements in performance over the zero-shot setting. However, trends become more variable as the number of ICL examples increases. In particular, we observe that increasing the number of ICL examples is not consistently correlated with stronger performance; in particular, all models exhibit substantial performance degradations that often dip below zero-shot performance. These results suggest that existing MLLMs may be unable to perform ICL tasks when provided with lengthy inputs consisting of multiple interleaved image-text pairs.
@@
Finally, in Figure 4 (Panel D), we evaluate MLLM performance stratified by image type. Each query image the SMMILE benchmark points to was annotated by an expert with one of the following categorical labels indicating modality: "X-Ray", "CT", "Staining", "Photograph", "ECG", "Ultrasound", "Fundus Photograph", "Text", "MRI", "Mammogram", "Illustration", "EEG", and "Other". We observe that MLLMs exhibit highly variable performance across the 13 considered modalities. No MLLM consistently achieves strong performance across all modalities, suggesting that the MLLMs are unable to consistently glean relevant information from provided in-context examples. We particularly note that when query images come from the MRI and Illustration modalities, all evaluated models fail to generate any correct answers. Modalities like "Text", "Mammogram", "Fundus Photograph", and "EEG" also prove to be challenging, with at least two MLLMs failing to generate any correct answers.
@@
In summary, our results demonstrate that SMMILE is a comprehensive and challenging benchmark for evaluating in-context learning abilities of MLLMs in medical settings. We hope that SMMILE can serve as a valuable resource for driving forward the development of future MLLMs.
@@ 4.1 Analyzing Example Quality
The role of in-context example quality on MLLM performance is not well-understood, and the high-quality nature of SMMILE provides a unique opportunity for addressing this question. Here, we create two perturbed versions of the SMMILE dataset as follows. (1) SMMILE-Random-Noise: For each sample in SMMILE, we add a random image-question-answer triplet from the dataset to the in-context example set. (2) SMMILE-Targeted-Noise: For each sample in SMMILE, we select an image-question-answer triplet from the dataset that shares the same specialty as the sample.
In Table 3, we report performance of 9 MLLMs across these perturbed variants of SMMILE. We observe that the inclusion of just one noisy sample in the in-context example list can impair performance, with most models exhibiting performance degradations on both SMMILE-Random-Noise (9.1% relative decrease from SMMILE on average) and SMMILE-Targeted-Noise (9.5% relative decrease from SMMILE on average). Targeted noise contributes to slightly lower performance than random noise on average, suggesting that even targeted, specialty-based selection of in-context examples can impair performance if the selected examples are not effective demonstrations of the task at hand. Importantly, the effects of noise are model-specific; the presence of noisy in-context examples affects each model in differing ways, leading to substantial changes in the final rankings. Our results demonstrate the critical need for high-quality, manually-curated benchmarks for evaluating in-context abilities of MLLMs in the medical setting, as the presence of noisy or irrelevant samples in the in-context example set can prevent developers from accurately understanding the capabilities of trained models.
Table 3: We create two perturbed verions of the SMMILE dataset (SMMILE-Random-Noise and SMMILE-Targeted-Noise) in order to evaluate the role of in-context example quality on MLLM performance. Here, we report performance across nine open-source MLLMs (ordered by model size) in the open-ended setting with LLM-as-a-Judge evaluation. The best result per row is bolded.

### v2

@@ 3.2 Benchmarking MLLMs with SMMILE
LLM-as-a-Judge Expert Rating EM MCQA
0-shot ICL 0-shot ICL ICL 0-shot ICL
Majority $26.30_{\scriptstyle{\pm 3.88}}$ $27.26_{\scriptstyle{\pm 4.03}}$ $24.15_{\scriptstyle{\pm 3.70}}$
Random $27.86_{\scriptstyle{\pm 4.43}}$ $23.16_{\scriptstyle{\pm 3.88}}$ $36.30_{\scriptstyle{\pm 4.66}}$
Text only $5.32_{\scriptstyle{\pm 2.27}}$ $16.53_{\scriptstyle{\pm 3.94}}$ $5.22_{\scriptstyle{\pm 1.70}}$ $38.62_{\scriptstyle{\pm 4.61}}$ $28.03_{\scriptstyle{\pm 4.28}}$
Claude 3.7 Sonnet $\mathbf{37.18}_{\scriptstyle{\pm 4.39}}$ $36.17_{\scriptstyle{\pm 4.44}}$ $\mathbf{39.19}$ $\mathbf{44.14}$ $2.63_{\scriptstyle{\pm 1.67}}$ $\mathbf{56.10}_{\scriptstyle{\pm 4.31}}$ $42.01_{\scriptstyle{\pm 4.83}}$
GPT-4o $32.56_{\scriptstyle{\pm 4.60}}$ $\mathbf{49.88}_{\scriptstyle{\pm 4.69}}$ $33.33$ $43.24$ $8.94_{\scriptstyle{\pm 2.54}}$ $49.74_{\scriptstyle{\pm 4.48}}$ $\mathbf{58.85}_{\scriptstyle{\pm 4.62}}$
Llama-3.2-Vision-90B $31.84_{\scriptstyle{\pm 4.38}}$ $40.66_{\scriptstyle{\pm 4.99}}$ $36.94$ $33.33$ $\mathbf{30.53}_{\scriptstyle{\pm 4.07}}$ $55.04_{\scriptstyle{\pm 4.93}}$ $30.30_{\scriptstyle{\pm 5.20}}$
LLaVA-v1.5-7B $14.61_{\scriptstyle{\pm 3.57}}$ $18.72_{\scriptstyle{\pm 3.40}}$ $17.12$ $21.62$ $16.37_{\scriptstyle{\pm 3.32}}$ $40.34_{\scriptstyle{\pm 5.35}}$ $22.30_{\scriptstyle{\pm 3.92}}$
LLaVA-v1.5-13B $19.58_{\scriptstyle{\pm 3.64}}$ $0.91_{\scriptstyle{\pm 3.49}} 22$.5222.52 $26.13$ $19.54_{\scriptstyle{\pm 3.95}}$ $38.96_{\scriptstyle{\pm 4.83}}$ $24.92_{\scriptstyle{\pm 4.25}}$
LLaVA-NeXT-7B $11.42_{\scriptstyle{\pm 3.04}}$ $23.66_{\scriptstyle{\pm 3.90}}$ $13.51$ $33.33$ $2.69_{\scriptstyle{\pm 1.31}}$ $38.11_{\scriptstyle{\pm 4.26}}$ $29.01_{\scriptstyle{\pm 4.05}}$
LLaVA-OneVision-0.5B $18.26_{\scriptstyle{\pm 3.93}}$ $21.63_{\scriptstyle{\pm 4.00}}$ $19.82$ $18.02$ $13.46_{\scriptstyle{\pm 3.11}}$ $44.03_{\scriptstyle{\pm 4.47}}$ $32.11_{\scriptstyle{\pm 4.44}}$
LLaVA-OneVision-7B $16.41_{\scriptstyle{\pm 3.65}}$ $24.25_{\scriptstyle{\pm 3.81}}$ $25.23$ $27.03$ $22.43_{\scriptstyle{\pm 4.33}}$ $40.15_{\scriptstyle{\pm 4.59}}$ $27.17_{\scriptstyle{\pm 4.28}}$
LLaVA-Med-7B∗∗ $21.65_{\scriptstyle{\pm 4.18}}$ $10.19_{\scriptstyle{\pm 3.06}}$ $22.52$ $10.81$ $_{\scriptstyle{\pm 0.00}} 0.00±0.00$$_{\scriptstyle{\pm 0.00}}$ $0.00_{\scriptstyle{\pm 0.00}}$
MedGemma-4B-Multimodal $27.73_{\scriptstyle{\pm 4.70}}$ $36.86_{\scriptstyle{\pm 4.81}}$ $27.03$ $32.43$ $12.14_{\scriptstyle{\pm 3.07}}$ $41.21_{\scriptstyle{\pm 5.15}}$ $40.67_{\scriptstyle{\pm 4.93}}$
MedVLM-R1 $25.20_{\scriptstyle{\pm 4.09}}$ $26.74_{\scriptstyle{\pm 4.44}}$ $25.23$ $24.32$ $15.26_{\scriptstyle{\pm 2.91}}$ $36.54_{\scriptstyle{\pm 4.63}}$ $26.22_{\scriptstyle{\pm 4.31}}$
Qwen2.5-VL-3B $27.62_{\scriptstyle{\pm 4.26}}$ $33.58_{\scriptstyle{\pm 4.09}}$ $23.42$ $23.42$ $26.18_{\scriptstyle{\pm 4.52}}$ $37.58_{\scriptstyle{\pm 4.41}}$ $27.35_{\scriptstyle{\pm 4.23}}$
Qwen2.5-VL-7B $17.90_{\scriptstyle{\pm 3.20}}$ $29.58_{\scriptstyle{\pm 4.63}}$ $11.71$ $27.03$ $22.45_{\scriptstyle{\pm 3.81}}$ $38.24_{\scriptstyle{\pm 4.75}}$ $45.01_{\scriptstyle{\pm 4.39}}$
Qwen2.5-VL-32B $25.27_{\scriptstyle{\pm 3.86}}$ $41.79_{\scriptstyle{\pm 4.73}}$ $21.62$ $32.43$ $31.84_{\scriptstyle{\pm 4.37}}$ $51.76_{\scriptstyle{\pm 4.46}}$ $49.97_{\scriptstyle{\pm 5.07}}$
Qwen2.5-VL-72B $29.90_{\scriptstyle{\pm 4.08}}$ $42.59_{\scriptstyle{\pm 4.55}}$ $26.13$ $31.52$ $15.71_{\scriptstyle{\pm 3.30}}$ $52.33_{\scriptstyle{\pm 4.58}}$ $54.71_{\scriptstyle{\pm 4.89}}$
@@
  • There are notable changes in model rankings. Unlike Table 1 where GPT-4o was dominant, Qwen2.5-VL-72B now takes the lead (53.8% LLM-as-a-Judge accuracy, 63.2% on ICL-MCQA).
  • Broader benefits from in-context learning are visible. Larger relative performance improvements are observed when in-context examples are presented to models, such as: LLaVA-v1.5-7B (99.4% relative improvement in LLM-as-a-Judge, 12.31% → 24.55%), Qwen2.5-VL-7B (94.4% relative improvement in LLM-as-a-Judge, 21.10% → 41.01%), and Qwen2.5-VL-3B (79.9% relative improvement in LLM-as-a-Judge, 17.53% → 31.54%). In the open-ended setting, all models (with the notable exception of LLaVA-Med-7B) demonstrate higher ICL performance than zero-shot performance when evaluated with LLM-as-a-Judge, with an average relative improvement of 44.7%.
@@
Majority - $17.70\pm\scriptstyle{1.19}$ - $17.59\pm\scriptstyle{1.19}$ - $16.65\pm\scriptstyle{1.13}$
Random - $25.35\pm\scriptstyle{1.34}$ - $25.35\pm\scriptstyle{1.32}$ - $33.77\pm\scriptstyle{1.46}$
@@
GPT-4o $\mathbf{38.41}\pm\scriptstyle{1.51}$ $46.45\pm\scriptstyle{1.54}$ $0.00\pm\scriptstyle{0.00}$ $7.79\pm\scriptstyle{0.81}$ $6.70\pm\scriptstyle{1.48} 55$$5.76\pm\scriptstyle{1.56}$
@@
LLaVA-Med-7B∗∗ $24.84\pm\scriptstyle{1.31}$ $4.62\pm\scriptstyle{0.64}$ $0.00\pm\scriptstyle{0.00}$ $0.19\pm\scriptstyle{0.14}$ $29\pm\scriptstyle{0.17} 0.0$$00\pm\scriptstyle{0.00}$
MedGemma-4B-Multimodal $24.72\pm\scriptstyle{1.35}$ $38.66\pm\scriptstyle{1.54}$ $0.00\pm\scriptstyle{0.00}$ $13.30\pm\scriptstyle{1.06}$ $0.36\pm\scriptstyle{1.55} 44$$4.78\pm\scriptstyle{1.52}$
@@ 3.3 Fine-Grained Analysis
We now perform a fine-grained breakdown of MLLM performance across the SMMILE benchmark. We specifically focus on five reproducible MLLMs for this analysis: LLaVA-OneVision-0.5B, LLaVA-Med-7B, LLaVA-v1.5-13B, Qwen2.5-VL-32B, and Qwen2.5-VL-72B. We first evaluate MLLM performance stratified by answer format. Each ground-truth answer in the SMMILE dataset was annotated by an expert with one of the following four categorical labels: "multiple choice", "free response", "binary (yes/no)", or "numerical". As shown in Figure 4 (Panel A), MLLMs display substantial variations in performance across the four categories, with all five evaluated models demonstrating the strongest performance on binary (yes/no) answers. Notably, we find that all evaluated models fail to correctly answer questions with numerical answers, which is a critical limitation since the ability to provide quantitative responses is vital for effective decision-making in medical settings. This finding is corroborated by analysis in Figure 4 (Panel B), which again finds that all evaluated MLLMs struggle when quantitative reasoning is necessary to answer a question.
@@
In Figure 4 (Panel C), we report the effect of the number of ICL examples on MLLM performance. For all evaluated models, providing two ICL examples leads to substantial improvements in performance over the zero-shot setting. However, trends become more variable as the number of ICL examples increases. In particular, we observe that increasing the number of ICL examples is not consistently correlated with stronger performance; in particular, all models exhibit substantial performance degradations that often dip below zero-shot performance. These results suggest that existing MLLMs may be unable to perform ICL tasks when provided with lengthy inputs consisting of multiple interleaved image-text pairs.
@@
Figure 4 (Panel D) shows that MLLMs exhibit highly variable performance across the 13 included imaging modalities. No MLLM achieves strong performance across all modalities, suggesting that the MLLMs are unable to consistently glean relevant information from provided in-context examples. When query images come from the MRI and illustration modalities, all evaluated models fail to generate any correct answers. The text, mammogram, fundus photograph, and EEG modalities also prove to be challenging, with at least two MLLMs failing to generate any correct answers.
@@
In summary, our results demonstrate that SMMILE is a comprehensive and challenging benchmark for evaluating in-context learning abilities of MLLMs in medical settings. We hope that SMMILE can serve as a valuable resource for driving forward the development of future MLLMs. Extended fine-grained analyses are provided in Appendix Sections E and F.
@@ 4.1 Analyzing Example Quality
The role of in-context example quality on MLLM performance is not well-understood, and the high-quality nature of SMMILE provides a unique opportunity for addressing this question. Here, we create two perturbed versions of the SMMILE dataset as follows. (1) SMMILE-Random-Noise: For each sample in SMMILE, we insert a random image-question-answer triplet from the dataset to the in-context example set. (2) SMMILE-Targeted-Noise: For each sample in SMMILE, we insert an image-question-answer triplet from the dataset that shares the same specialty as the sample.
In Table 3, we report performance of 9 MLLMs across these perturbed variants of SMMILE. We observe that the inclusion of just one noisy sample in the in-context example list can impair performance, with most models exhibiting performance degradations on both SMMILE-Random-Noise (9.1% relative decrease from SMMILE on average) and SMMILE-Targeted-Noise (9.5% relative decrease from SMMILE on average). Targeted noise contributes to slightly lower performance than random noise on average, suggesting that even targeted, specialty-based selection of in-context examples can impair performance if the selected examples are not effective demonstrations of the task at hand. Importantly, the effects of noise are model-specific; the presence of noisy in-context examples affects each model in differing ways, leading to substantial changes in the final rankings. Our results demonstrate the critical need for high-quality, manually-curated benchmarks for evaluating in-context abilities of MLLMs in the medical setting, as the presence of noisy or irrelevant samples in the in-context example set can prevent developers from accurately understanding model capabilities.
Table 3: We create two perturbed versions of the SMMILE dataset (SMMILE-Random-Noise and SMMILE-Targeted-Noise) in order to evaluate the role of in-context example quality on MLLM performance. Here, we report performance across nine open-source MLLMs (ordered by model size) in the open-ended setting with LLM-as-a-Judge evaluation. The best result per row is bolded.